A set of mid term examination grades in an undergraduate introductory statistics course at a large public university (i.e., there were lots of students) was found to be normally distributed with a mean of 73 and a standard deviation of 8.
1. What is the probability of getting between a 60 and 70 on this exam?

Answer by stanbon(75887)

normally distributed with a mean of 73 and a standard deviation of 8.
1. What is the probability of getting between a 60 and 70 on this exam?
----
z(60) = (60-73)/8 = -13/8 = -1.6250
---
z(70) = (70-73)/8 = -3/8 = -0.375
---
Using my TI-84 I get:
P(60< x <70) = P(-1.625< z <-0.375)
= normalcdf(-1.625,-0.375)
= 0.3017
=========================================
Cheers,
Stan H.
==========